CAT Chat Minutes

November 14, 1997

 

Issues that require action and/or follow-up
1. The unavailability of the linac during the last 24 hrs. has resulted in a run with 20 mA stored current. The general consensus from the users was that they were pleased with being able to use the 20 mA beam (see operational study below).

2. Regarding the undulator control, our screen indicated operator mode, but when we accidentally triggered an undulator scan, it began moving and the screen changed to User Mode. Should this have happened?

No it should not have happened. We will be looking into this.

3. Is it possible to put in a PV identifying the software version you are using for the undulator gap?

We have established revision control and will look into providing this data.

4. Regarding the calibration of undulators, we had mapped it out during the last cycle. We are now showing a significant change. It is now at 120 eV. What was changed?

We do not know what is causing this. Only a few CATs experienced this. No changes were made. We are looking into it and hope to have more information on this by the next run. What you are seeing (the 120 eV) is what we would like you to see.

5. PNC has lost their tricycle. If anyone sees a orange tricycle with PNC painted on it, please contact S. Heald at x 9795.

6. The APS seems to be a very dry place. The reason we have noticed this is that the birds are trying to get to our monochromator water supply. We have put up an alternate bird bath, but were wondering if something can be done about the birds on the Experiment Hall floor?

We will take action on this immediately.

7. There seems to be a problem with gas cylinders in the truck lock. They are not being picked up when they are empty. They are not AGA cylinders.

The problem is that CARS CAT has contracted with a different vender, and that vender is not picking up the empty cylinders. F. Bellinger will follow up with the CAT.

8. The door in the 432 LOM is working very well. Thank you.

Information/Follow-up Items
APS Operational Status Update:  We have had some linac problems. During the last maintenance period, some rf components were replaced. These components were mechanically superior, but electrically inferior, to the components that they replaced. The plan is to go back to the old components and then back to the vender with the new components, to get them checked out. We expect the linac to be operational tonight and to have injection tomorrow morning at 06:30.

PV Gateway Problems:  During the Monday ASD beam studies time, the PV gateway was tested in an attempt to fix the problems related to rebooting of the gateway due to IOC reboots. This problem was traced to IOC access security issues, which will be corrected by next month. In the course of the testing, it was detected that the running version of the gateway had been compiled on a different machine and an incorrect version of the ethernet driver had been installed, so a system patch was installed. On Tuesday morning, at the start of User Operations, it was noticed that the gateway on RHEA was not running. When the gateway was restarted, a problem with the access security file was noted. While the cause was being investigated, a backup copy was loaded. This placed the ID control in user mode, but access was still denied. By noon, the correct access security file and the correct version of the gateway software were loaded and user control of devices was reestablished. The floor coordinator control was still denied, and this problem was traced to the fact that the FCs access the gateway through HYDRA. On HYDRA, the gateway program ran under a different user name, which prevented FC access. During the shutdown, the FC access will be moved to RHEA as well.

ID Gap Closure Problem: At the end of User Operation runs preceding an ASD studies period, operating ID gap information is recorded by the BLOPS group. During SR studies, in testing operating modes and beam effects, the ID gaps are controlled by the machine physicist. Because the shutters are closed, the gaps may be closed below the normal operating limits. At the end of the studies period, the previously recorded operating gap information is reloaded.

On Tuesday, November 11, an incorrect sequence was used in restoring the gap information database and some of the devices were left fully closed. Procedures are being corrected to ensure that such occurrences will be avoided in the future. For the long term, we will investigate methods for incorporating protection directly into the ID controls and minimizing reliance on procedures.

Other
We are reminding all users that they have access to a PV that allows them determine a lower operational limit for the ID gap and have it set into the ID control software by the Floor Coordinator. These values are changed only by the Floor Coordinators under instruction from the beamline staff. The ID control program will not allow the devices to be operated at gaps below this set point. If there is risk to your beamline equipment from closing the ID gap, we strongly urge you to make use of this feature.

To remind you - if you are not using your device for a period longer than four hours, the FCs have been instructed to open it. Always keep your device open when you are not using it. We will open it if it is not open already.

We will be running the low beta lattice operating mode during machine studies on Monday 11-17-97 and again on Tuesday 11-18-97. This lattice has low beta value only in Sector 3, with the rest of the sectors having the standard beta value. If the tests are trouble free, we will continue this lattice during next week's operation. A final decision will be made on Tuesday next week, and you will be informed on CAT Net. We will also be doing shielding verification at 25 mA on Tuesday.